# http://dreamscene360.github.io/

### v1

@@ Abstract
The increasing demand for virtual reality applications has highlighted the significance of crafting immersive 3D assets. We present a text-to-3D ${}^{\circ}$ scene generation pipeline that facilitates the creation of comprehensive ${}^{\circ}$ scenes for in-the-wild environments in a matter of minutes. Our approach utilizes the generative power of a 2D diffusion model and prompt self-refinement to create a high-quality and globally coherent panoramic image. This image acts as a preliminary “flat” (2D) scene representation. Subsequently, it is lifted into 3D Gaussians, employing splatting techniques to enable real-time exploration. To produce consistent 3D geometry, our pipeline constructs a spatially coherent structure by aligning the 2D monocular depth into a globally optimized point cloud. This point cloud serves as the initial state for the centroids of 3D Gaussians. In order to address invisible issues inherent in single-view inputs, we impose semantic and geometric constraints on both synthesized and input camera views as regularizations. These guide the optimization of Gaussians, aiding in the reconstruction of unseen regions. In summary, our method offers a globally consistent 3D scene within a ${}^{\circ}$ perspective, providing an enhanced immersive experience over existing techniques. Project website at: http://dreamscene360.github.io/.
Figure 1: DreamScene360. We introduce a 3D scene generation pipeline that creates immersive scenes with full ${}^{\circ}$ coverage from text prompts of any level of specificity.
11footnotetext: Equal contribution.
@@ 1 Introduction
Moving beyond 2D, the generation of 3D content, particularly 3D scenes, is constrained by the limited availability of annotated 3D image-text data pairs. Consequently, efforts in 3D content creation often rely on leveraging large-scale 2D models. This line of approach facilitates the creation of 3D scenes through a time-consuming distillation process. An example of this is DreamFusion [45], which seeks to distill the object-wise 2D priors from diffusion models into a 3D neural radiance field (NeRF) [33]. However, these approaches often suffer from low rendering quality, primarily due to the multi-view inconsistency of 2D models, and struggle to extend to scene-scale 3D structure with fine details texture creation, particularly for outdoor scenes [45] with outward-facing viewpoints and unbounded scene scale. Another avenue of 3D generation draws insights from explicit representations, such as point clouds and meshes, as demonstrated in LucidDreamer [7] and Text2Room [20]. These methods attempt to bridge the gap between 2D and 3D generation by initializing with an explicit 3D representation, and then progressively expanding the learned 3D representation to encompass a broader field-of-view. However, the progressive optimization frameworks leveraged by these methods struggle to inpaint substantial missing areas, especially when targeting ${}^{\circ}$ scenes under unconstrained conditions, resulting in notably distorted and disjointed structures. Moreover, the issue of prompt engineering in text-to-image generation [50, 51], becomes more pronounced in text-to-3D generation frameworks [7, 45, 1] that rely on either time-consuming score distillation or complex, multi-step progressive inpainting during the scene generation process, leading to a considerable trial-and-error effort to achieve the desired 3D scene.
@@
To address the above challenges in creating a holistic ${}^{\circ}$ text-to-3D scene generation pipeline, we introduce DreamScene360. Our method initially leverages the generative capabilities of text-to-panorama diffusion models [60] to produce omnidirectional ${}^{\circ}$ panoramas providing a comprehensive representation of the scene. A self-refining mechanism is used to enhance the image to alleviate prompt engineering, where GPT-4V is integrated to improve the visual quality and the text-image alignment through iterative quality assessment and prompt revision. While the generated panorama images overcome the view consistency issue across different viewpoints, they still lack depth information and any layout priors in unconstrained settings, and contains partial observations due to their single-view nature. To address this, our approach involves initializing scale-consistent scene geometry by employing a pretrained monocular depth estimator alongside an optimizable geometric field, facilitating deformable alignment for each perspective-projected pixel. The gaps, stemming from single-view observations, can be filled by deforming the Gaussians to the unseen regions by creating a set of pseudo-views with a synthesized multi-view effect and the distillation of pseudo geometric and semantic constraints from 2D models (DPT [48] and DINO-ViT [42]) to guide the deformation process to alleviate artifacts.
@@ 2 Related Works
The generation of 2D assets allows for incredible creative liberty, and the use of large-scale learning based priors for content generation. Generative Adversarial Networks [15] were the original state of the art for image generation. Variants such as StyleGAN [24] showed the ability for fine-grained control of attributes such as expression through manipulation of the latent representations as well. More recently, denoising diffusion models [9, 18, 55] have been the new state of the art for generative models. Text-guided image diffusion models [52] have shown the capability of generating high-quality images that are faithful to the conditioning text prompts. Subsequent work has made the generation process more efficient by performing denoising in the latent space [50], and by speeding up the denoising process [53, 32, 14]. Techniques such as classifier free guidance [19] have significantly improved faithfulness to text prompts. Additional control of generation has also been shown to be possible, through auxiliary inputs such as layout [71], pose [70] and depth maps [5]. More recently, text-to-image diffusion models have been finetuned to generate structured images, such as panoramas [60]. In this work, we utilize text-to-panorama generation as structured guidance for text to ${}^{\circ}$ scene generation.
@@
In recent times, text-to-3D scene generation has been reinvigorated with the advent of 2D diffusion models and guidance techniques. Text-to-image diffusion models, along with techniques such as classifier free guidance, have been found to provide strong priors to guide text-to-3D generation methods [45, 62, 61]. These methods are generally more frequently used than direct text-to-3D diffusion models [22, 23]. More recent works [29, 13, 58] successfully generate multi-object compositional 3D scenes. On the other hand, a second class of methods use auxiliary inputs such as layouts [44]. The least constrained text-to-3D methods do not rely on any auxiliary input with the only input being the text prompt describing the 3D scene [20, 69, 43, 11, 31, 67, 7]. Our work requires a text prompt input; however, unlike prior work, we propose using panoramic images as an intermediate input for globally consistent scenes.
@@
3D scene representations include a wide variety of techniques, including point clouds [4], volumetric representations [30, 40], and meshes [68]. More recently, however, learning-based scene representations have gained prominence. Implicit representations such as neural radiance fields [34] have shown high quality rendering and novel view synthesis capability. Additional learning-based representations, that span both explicit [12] and mixed [38] have been proposed to speed up learning. Most recently, 3D Gaussian splatting [25] has enabled fast learning and rendering of high quality radiance fields using a Gaussian kernel-based explicit 3D representation. Since then, several works have emerged enabling sparse view [73, 64] and compressed [41, 10, 39, 27, 37] 3D Gaussian representations, as well as representation of multidimensional feature fields [72]. A variation involves learning a Gaussian-based radiance field representation using panoramic images as input [2]. In this work, we propose a method for text to ${}^{\circ}$ 3D scene generation, by using panorama images as an intermediate representation.
@@ 3 Methods
In this section, we detail the proposed DreamScene360 architecture (Fig.2). DreamScene360 initially generates a ${}^{\circ}$ panorama utilizing a self-refinement process, which ensures robust generation and aligns the image with the text semantically (Sec. 3.1). The transformation from a flat 2D image to a 3D model begins with the initialization from a panoramic geometric field (Sec. 3.2), followed by employing semantic alignment and geometric correspondences as regularizations for the deformation of the 3D Gaussians (Sec. 3.3).
Figure 2: Overall Architecture. Beginning with a concise text prompt, we employ a diffusion model to generate a ${}^{\circ}$ panoramic image. A self-refinement process is employed to produce the optimal 2D candidate panorama. Subsequently, a 3D geometric field is utilized to initialize the Panoramic 3D Gaussians. Throughout this process, both semantic and geometric correspondences are employed as guiding principles for the optimization of the Gaussians, aiming to address and fill the gaps resulting from the single-view input.
3.1 Text to ${}^{\circ}$ Panoramas with Self-Refinement
Panoramic images provide an overview of the entire scene in one image, which is essential for generating ${}^{\circ}$ 3D scenes with global consistency.
${}^{\circ}$ Panoramic Image Generation.
@@ 3.1 Text to ${}^{\circ}$ Panoramas with Self-Refinement
Similar to StitchDiffusion [60], we employ a stitch method in the generation process for synthesizing seamless ${}^{\circ}$ panoramic images. Trained on a curated paired image-text dataset containing ${}^{\circ}$ panoramas, a customized LoRA [21] module is incorporated into the MultiDiffusion [3] process. At each denoising timestamp, we not only diffuse at the original resolution $H\times(2H+2W)$ via MultiDiffuser $\Phi$, but also stitch the leftmost and rightmost image regions ($H\times W$ each) and diffuse the concatenated patch as well to ensure consistency at the border regions. Finally, we consider the center cropped region of $H\times 2H$ as the final ${}^{\circ}$ panoramic image.
@@
Our 3D scene generation framework, therefore, enjoys a user-friendly self-improvement process, without the need for troublesome prompt engineering for users as in previous methods [45, 62, 7], but is able to obtain a high-quality, visually pleasing, text-aligned, and ${}^{\circ}$ consistent panoramic images that can be later converted into an immersive 3D scene via Panoramic Gaussian Splatting, which is detailed in the next sections.
@@ 3.2 Lifting in-the-wild Panorama to 360 Scene
Transforming a single image, specifically an in-the-wild ${}^{\circ}$ panoramic image, into a 3D model poses significant challenges due to inadequate observational data to regularize the optimization process, such as those required in 3D Gaussian Splatting (see Fig. 7). Rather than beginning with a sparse point cloud (3DGS), we initialize with a dense point cloud utilizing pixel-wise depth information from the panoramic image of resolution H $\times$ W, which are then refined towards more precise spatial configurations, ensuring the creation of globally consistent 3D representations, robust to viewpoint changes.
@@
Previous studies in deformable depth alignment [16, 59, 49] and pose-free novel view synthesis [6] have explored aligning scales across multiple view depth maps derived from monocular depth estimations. In this context, we utilize a learnable global geometric field (MLPs), supplemented by per-view scale and per-pixel shift parameters: $\{(\alpha_{i}\in\mathbb{R},\boldsymbol{\beta_{i}}\in\mathbb{R}^{H\times W}\}_{i=1}^{N}$. We define the view direction ($\boldsymbol{v}$) for all pixels on the perspective images deterministically. The parameters of MLPs $\Theta$ are initialized with an input dimension of three and an output dimension of one. The parameters $\{(\alpha_{i}\in\mathbb{R}\}_{i=1}^{N}$ are initialized to ones, and $\{(\boldsymbol{\beta_{i}}\in\mathbb{R}^{H\times W}\}_{i=1}^{N}$ to zeros. With these optimizable parameters, we define our optimization goal as follows:
$\min_{\alpha,\beta,\Theta}\bigg{\{}||\alpha\cdot\boldsymbol{D}^{\text{Mono}}+\boldsymbol{\beta}-\operatorname{MLPs}(\boldsymbol{v};\Theta)||_{2}^{2}$ (2)
$\quad+\lambda_{\text{TV}}\mathcal{L}_{\text{TV}}(\boldsymbol{\beta})+\lambda_{\alpha}||\text{softplus}(\alpha)-1||^{2}\bigg{\}}$
@@
Here, $\lambda_{\text{TV}}$ and $\lambda_{\alpha}$ are regularization coefficients to balance the loss weight. We set FoV as ${}^{\circ}$ during the optimization. We will use the predicted depth from the MLPs for the subsequent Gaussian optimization, as it can provide the depth of any view direction.
@@ 3.3 Optimizing Monocular Panoramic 3D Gaussians
Previous research has underscored the significance of capturing the appearance of objects/scenes to establish robust feature correspondences across different views [17, 28], which is crucial for tasks like co-segmentation and 2D pre-training. Inspired by these studies, we aim to establish a connection in visual feature correspondence between training views and synthetically generated virtual views. This is achieved by enforcing feature-level similarity, guiding the 3D Gaussians to fill the geometric gaps in the invisible regions effectively. To elaborate, we generate a perspective image $\boldsymbol{I}_{i}$ from the panoramic 3D Gaussians at a specific training camera viewpoint $\boldsymbol{P}_{i}$ and create another synthesized image $\boldsymbol{I}^{\prime}_{i}$ using Eq 4 that is proximate to $\boldsymbol{P}_{i}$. We employ the [CLS] token from the pre-trained DINO-ViT [42] model to encapsulate compact semantic features through the equation:
[…]
where $\operatorname{Cos}$ denotes the cosine similarity. This approach is predicted on a similar assumption to that of Tumanyan et al. [57], assuming that the [CLS] token from a self-supervised, pre-trained Vision Transformer (ViT) can capture the high-level semantic attributes of an image. Such a mechanism is instrumental in identifying and leveraging similarities between two adjacent rendered images, facilitating our end-to-end 3D-GS optimization process.
@@
While leveraging appearance cues is valuable, relying solely on them can lead to spatial inconsistencies. This is because models like DINO-ViT might prioritize capturing semantic features over maintaining geometric coherence, potentially resulting in artifacts such as floaters. To mitigate this, we introduce a geometric regularization strategy designed to penalize discontinuities between pixels that exhibit inaccurate depth relationships. We employ a monocular depth estimator DPT [48] for this purpose [8]. Although it may not provide globally accurate depth, it allows us to ascertain the relative spatial relationships between pixels:
[…]
@@
3.4 Optimization
[…]
Figure 4: Visual Comparisons. We showcase ${}^{\circ}$ 3D scene generation. In each row, from left to right, displays novel views as the camera undergoes clockwise rotation in yaw, accompanied by slight random rotations in pitch and random translations. LucidDreamer [7] hallucinates novel views from a conditioned image (indicated by a red bounding box) but lacks global semantic, stylized, and geometric consistency. In contract, our method provides complete ${}^{\circ}$ coverage without any blind spots (black areas in baseline results), and show globally consistent semantics.
4 Experiments
4.1 Experiment Setting
Implementation Details.
We first generate panoramas at the resolution of $512\times 1024$, following the setting of [60]. Then, we upsample the panorama image to the size of $1024\times 2048$ via bilinear interpolation, in order to produce a dense enough point cloud by projecting per-pixel panoramic depth values into 3D space along the per-pixel camera ray directions, where the directions can be obtained according to the coordinate transformation in spherical panoramic imaging [49]. Since the panoramic depth map is predicted from the optimized geometric field, the values in local patches are guaranteed to be consistent and smooth, thereby resulting in point cloud in 3D that accurately captures the smooth geometry at the objects’ surfaces. We call this process as lifting. For unbounded scenes, points representing distant objects (like the sky in outdoor settings) do not have infinite coordinates due to the finite depth values from the 2D panoramic depth map. This setup enables effective point cloud initialization within a controllable, bounded area, eliminating redundant points in empty spaces. We leverage this point cloud as the foundation for 3D Gaussian Splatting (3DGS) initialization. To circumvent the common floater issue in 3DGS rendering, we disable the densification process, enhancing the overall quality and consistency of the rendered scenes. We start the 3-stage progressive camera perturbation from 5400 iterations during the training, the interval of each stage is adjustable, and we set it to 1200 iterations in practice. With the same observation in [25], most of the scenes at 7000 iterations can already achieve very high quality.
Baseline Methods.
Our work tackles the challenging problem of unconstrained ${}^{\circ}$ generation, including both indoor and outdoor scenarios. However, the works utilizing a bounded NeRF representation using score distillation do not work very well in this case. Thus, the comparisons are conducted between DreamScene360 (ours) and the state-of-the-art LucidDreamer [7]. We use the open-source codebase of LucidDreamer, which starts from a single image and a text prompt. The framework constructs a global point cloud by progressive inpainting into ${}^{\circ}$ views and then distills a set of 3D Gaussians. In our experiments, we set the input image for the LucidDreamer to be generated from Stable Diffusion [50] v1.5 using the input text for a fair comparison.
Metrics.
4.2 Main Results
${}^{\circ}$ Scene Generation.
Comparisons with Baseline Methods.
We show visual comparisons against LucidDreamer [7] in Fig. 4. LucidDreamer involves progressive inpainting to find agreement between multiple synthesized images and tries to fuse a point cloud that can be later distilled into 3D Gaussians. Their pipeline tends to produce recurrent results because each patch is inpainted separately given the same text prompt. This is cumbersome especially when generating scenes with complex structures, for example, the street view. In comparison, our method delivers consistent results thanks to the intermediate panorama as a global 2D representation. We provide quantitative comparisons in Tab. 1, where the camera undergoes clockwise rotation in yaw, accompanied by slight random rotations in pitch and random translations to capture 4 novel views roughly representing front, back, left, and right, along with two views of up and down by ${}^{\circ}$ rotation in pitch, to mimic the exploration in an immersive ${}^{\circ}$ panoramic 3D scene.
For the bedroom text prompt, LucidDreamer begins with a view featuring a modern style bedroom with one bed but hallucinates the novel views into multiple bedrooms, while the style of the imagined bedrooms diverges from that of the starting view. In the case of the Yosemite text prompt, LucidDreamer merely replicates the waterfall seen in the initial view throughout. In conclusion, our results demonstrate global semantic, stylized, and geometric consistency, offering complete ${}^{\circ}$ coverage without any blind spots.

### v2

# %****␣main.tex␣Line␣100␣****http://dreamscene360.github.io/

@@ Abstract
The increasing demand for virtual reality applications has highlighted the significance of crafting immersive 3D assets. We present a text-to-3D 360 scene generation pipeline that facilitates the creation of comprehensive 360 scenes for in-the-wild environments in a matter of minutes. Our approach utilizes the generative power of a 2D diffusion model and prompt self-refinement to create a high-quality and globally coherent panoramic image. This image acts as a preliminary “flat” (2D) scene representation. Subsequently, it is lifted into 3D Gaussians, employing splatting techniques to enable real-time exploration. To produce consistent 3D geometry, our pipeline constructs a spatially coherent structure by aligning the 2D monocular depth into a globally optimized point cloud. This point cloud serves as the initial state for the centroids of 3D Gaussians. In order to address invisible issues inherent in single-view inputs, we impose semantic and geometric constraints on both synthesized and input camera views as regularizations. These guide the optimization of Gaussians, aiding in the reconstruction of unseen regions. In summary, our method offers a globally consistent 3D scene within a 360 perspective, providing an enhanced immersive experience over existing techniques. Project website at: http://dreamscene360.github.io/.
Figure 1: DreamScene360. We introduce a 3D scene generation pipeline that creates immersive scenes with full 360 coverage from text prompts of any level of specificity.
@@ 1 Introduction
Moving beyond 2D, the generation of 3D content, particularly 3D scenes, is constrained by the limited availability of annotated 3D image-text data pairs. Consequently, efforts in 3D content creation often rely on leveraging large-scale 2D models. This line of approach facilitates the creation of 3D scenes through a time-consuming distillation process. An example of this is DreamFusion [46], which seeks to distill the object-wise 2D priors from diffusion models into a 3D neural radiance field (NeRF) [34]. However, these approaches often suffer from low rendering quality, primarily due to the multi-view inconsistency of 2D models, and struggle to extend to scene-scale 3D structure with fine details texture creation, particularly for outdoor scenes [46] with outward-facing viewpoints and unbounded scene scale. Another avenue of 3D generation draws insights from explicit representations, such as point clouds and meshes, as demonstrated in LucidDreamer [7] and Text2Room [20]. These methods attempt to bridge the gap between 2D and 3D generation by initializing with an explicit 3D representation, and then progressively expanding the learned 3D representation to encompass a broader field-of-view. However, the progressive optimization frameworks leveraged by these methods struggle to inpaint substantial missing areas, especially when targeting 360 scenes under unconstrained conditions, resulting in notably distorted and disjointed structures. Moreover, the issue of prompt engineering in text-to-image generation [51, 52], becomes more pronounced in text-to-3D generation frameworks [7, 46, 1] that rely on either time-consuming score distillation or complex, multi-step progressive inpainting during the scene generation process, leading to a considerable trial-and-error effort to achieve the desired 3D scene.
@@
To address the above challenges in creating a holistic 360 text-to-3D scene generation pipeline, we introduce DreamScene360. Our method initially leverages the generative capabilities of text-to-panorama diffusion models [62] to produce omnidirectional 360 panoramas providing a comprehensive representation of the scene. A self-refining mechanism is used to enhance the image to alleviate prompt engineering, where GPT-4V is integrated to improve the visual quality and the text-image alignment through iterative quality assessment and prompt revision. While the generated panorama images overcome the view consistency issue across different viewpoints, they still lack depth information and any layout priors in unconstrained settings, and contains partial observations due to their single-view nature. To address this, our approach involves initializing scale-consistent scene geometry by employing a pretrained monocular depth estimator alongside an optimizable geometric field, facilitating deformable alignment for each perspective-projected pixel. The gaps, stemming from single-view observations, can be filled by deforming the Gaussians to the unseen regions by creating a set of pseudo-views with a synthesized multi-view effect and the distillation of pseudo geometric and semantic constraints from 2D models (DPT [49] and DINOv2 [43]) to guide the deformation process to alleviate artifacts.
@@ 2 Related Works
The generation of 2D assets allows for incredible creative liberty, and the use of large-scale learning based priors for content generation. Generative Adversarial Networks [15] were the original state of the art for image generation. Variants such as StyleGAN [24] showed the ability for fine-grained control of attributes such as expression through manipulation of the latent representations as well. More recently, denoising diffusion models [9, 18, 56] have been the new state of the art for generative models. Text-guided image diffusion models [53] have shown the capability of generating high-quality images that are faithful to the conditioning text prompts. Subsequent work has made the generation process more efficient by performing denoising in the latent space [51], and by speeding up the denoising process [54, 33, 14]. Techniques such as classifier free guidance [19] have significantly improved faithfulness to text prompts. Additional control of generation has also been shown to be possible, through auxiliary inputs such as layout [74], pose [73] and depth maps [5]. More recently, text-to-image diffusion models have been finetuned to generate structured images, such as panoramas [62]. In this work, we utilize text-to-panorama generation as structured guidance for text to 360 scene generation.
@@
In recent times, text-to-3D scene generation has been reinvigorated with the advent of 2D diffusion models and guidance techniques. Text-to-image diffusion models, along with techniques such as classifier free guidance, have been found to provide strong priors to guide 3D generation methods [46, 64, 63, 57, 68]. These methods are generally more frequently used than direct text-to-3D diffusion models [22, 23]. More recent works [30, 13, 60] successfully generate multi-object compositional 3D scenes. On the other hand, a second class of methods use auxiliary inputs such as layouts [45]. The least constrained text-to-3D methods do not rely on any auxiliary input with the only input being the text prompt describing the 3D scene [20, 72, 44, 11, 32, 70, 7]. Our work requires a text prompt input; however, unlike prior work, we propose using panoramic images as an intermediate input for globally consistent scenes.
@@
3D scene representations include a wide variety of techniques, including point clouds [4], volumetric representations [31, 41], and meshes [71]. More recently, however, learning-based scene representations have gained prominence. Implicit representations such as neural radiance fields [35] have shown high quality rendering and novel view synthesis capability. Additional learning-based representations, that span both explicit [12] and mixed [39] have been proposed to speed up learning. Most recently, 3D Gaussian splatting [25] has enabled fast learning and rendering of high quality radiance fields using a Gaussian kernel-based explicit 3D representation. Since then, several works have emerged enabling sparse view [76, 66] and compressed [42, 10, 40, 27, 38] 3D Gaussian representations, as well as representation of multidimensional feature fields [75]. A variation involves learning a Gaussian-based radiance field representation using panoramic images as input [2]. In this work, we propose a method for text to 360 3D scene generation, by using panorama images as an intermediate representation.
@@ 3 Methods
In this section, we detail the proposed DreamScene360 architecture (Fig.2). DreamScene360 initially generates a 360 panorama utilizing a self-refinement process, which ensures robust generation and aligns the image with the text semantically (Sec. 3.1). The transformation from a flat 2D image to a 3D model begins with the initialization from a panoramic geometric field (Sec. 3.2), followed by employing semantic alignment and geometric correspondences as regularizations for the deformation of the 3D Gaussians (Sec. 3.3).
Figure 2: Overall Architecture. Beginning with a concise text prompt, we employ a diffusion model to generate a 360 panoramic image. A self-refinement process is employed to produce the optimal 2D candidate panorama. Subsequently, a 3D geometric field is utilized to initialize the Panoramic 3D Gaussians. Throughout this process, both semantic and geometric correspondences are employed as guiding principles for the optimization of the Gaussians, aiming to address and fill the gaps resulting from the single-view input.
3.1 Text to 360 Panoramas with Self-Refinement
Panoramic images provide an overview of the entire scene in one image, which is essential for generating 360 3D scenes with global consistency.
360 Panoramic Image Generation.
@@ 3.1 Text to 360 Panoramas with Self-Refinement
We use StitchDiffusion [62] as the pretrained 2D diffusion model, where a stitch method is employed in the generation process for synthesizing seamless 360 panoramic images. Trained on a curated paired image-text dataset containing 360 panoramas, a customized LoRA [21] module is incorporated into the MultiDiffusion [3] process. At each denoising timestamp, we not only diffuse at the original resolution $H\times(2H+2W)$ via MultiDiffuser $\Phi$, but also stitch the leftmost and rightmost image regions ($H\times W$ each) and diffuse the concatenated patch as well to ensure consistency at the border regions. Finally, we consider the center cropped region of $H\times 2H$ as the final 360 panoramic image. It is worth noting that Dreamscene360 is versatile in practice and can also adapt to other text-to-panorama diffusion models.
@@
Our 3D scene generation framework, therefore, enjoys a user-friendly self-improvement process, without the need for troublesome prompt engineering for users as in previous methods [46, 64, 7], but is able to obtain a high-quality, visually pleasing, text-aligned, and 360 consistent panoramic images that can be later converted into an immersive 3D scene via Panoramic Gaussian Splatting, which is detailed in the next sections.
@@ 3.2 Lifting in-the-wild Panorama to 360 Scene
Transforming a single image, specifically an in-the-wild 360 panoramic image, into a 3D model poses significant challenges due to inadequate observational data to regularize the optimization process, such as those required in 3D Gaussian Splatting (see Fig. 7). Rather than beginning with a sparse point cloud (3DGS), we initialize with a dense point cloud utilizing pixel-wise depth information from the panoramic image of resolution H $\times$ W, which are then refined towards more precise spatial configurations, ensuring the creation of globally consistent 3D representations, robust to viewpoint changes.
@@
Previous studies in deformable depth alignment [16, 61, 50] and pose-free novel view synthesis [6] have explored aligning scales across multiple view depth maps derived from monocular depth estimations. In this context, we utilize a learnable global geometric field (MLPs), inspired by [61, 50], supplemented by per-view scale and per-pixel shift parameters: $\{(\alpha_{i}\in\mathbb{R},\boldsymbol{\beta_{i}}\in\mathbb{R}^{H\times W}\}_{i=1}^{N}$. We define the view direction ($\boldsymbol{v}$) for all pixels on the perspective images deterministically. The parameters of MLPs $\Theta$ are initialized with an input dimension of three and an output dimension of one. The parameters $\{(\alpha_{i}\in\mathbb{R}\}_{i=1}^{N}$ are initialized to ones, and $\{(\boldsymbol{\beta_{i}}\in\mathbb{R}^{H\times W}\}_{i=1}^{N}$ to zeros. With these optimizable parameters, we define our optimization goal as follows:
$\min_{\alpha,\beta,\Theta}\bigg{\{}||\alpha\cdot\boldsymbol{D}^{\text{Mono}}+\boldsymbol{\beta}-\operatorname{MLPs}(\boldsymbol{v};\Theta)||_{2}^{2}+\lambda_{\text{TV}}\mathcal{L}_{\text{TV}}(\boldsymbol{\beta})+\lambda_{\alpha}||\gamma(\alpha)-1||^{2}\bigg{\}}$ (2)
@@
Here, $\lambda_{\text{TV}}$ and $\lambda_{\alpha}$ are regularization coefficients to balance the loss weight, $\gamma(\cdot)$ is the softplus function. We set FoV as 80 during the optimization and use the predicted depth from the MLPs for the subsequent Gaussian optimization, as it can provide the depth of any view direction.
@@ 3.3 Optimizing Monocular Panoramic 3D Gaussians
Previous research has underscored the significance of capturing the appearance of objects/scenes to establish robust feature correspondences across different views [17, 29], which is crucial for tasks like co-segmentation and 2D pre-training. Inspired by these studies, we aim to establish a connection in visual feature correspondence between training views and synthetically generated virtual views. This is achieved by enforcing feature-level similarity, guiding the 3D Gaussians to fill the geometric gaps in the invisible regions effectively. To elaborate, we generate a perspective image $\boldsymbol{I}_{i}$ from the panoramic 3D Gaussians at a specific training camera viewpoint $\boldsymbol{P}_{i}$ and create another synthesized image $\boldsymbol{I}^{\prime}_{i}$ using Eq 4 that is proximate to $\boldsymbol{P}_{i}$. We employ the [CLS] token from the pre-trained DINOv2 [43] model to encapsulate compact semantic features through the equation:
$\mathcal{L}_{\textit{sem}}=1-\operatorname{Cos}(\text{[CLS]}(I_{i}),\text{[CLS]}(I^{\prime}_{i})),$ (5)
where $\operatorname{Cos}$ denotes the cosine similarity. This approach is inspired by [59, 67], assuming that the [CLS] token from a self-supervised, pre-trained Vision Transformer (ViT) can capture the high-level semantic attributes of an image. Such a mechanism is instrumental in identifying and leveraging similarities between two adjacent rendered images, facilitating our end-to-end 3DGS optimization process.
@@
While leveraging appearance cues is valuable, relying solely on them can lead to spatial inconsistencies. This is because models like DINOv2 might prioritize capturing semantic features over maintaining geometric coherence, potentially resulting in artifacts such as floaters. To mitigate this, we introduce a geometric regularization strategy designed to penalize discontinuities between pixels that exhibit inaccurate depth relationships. We employ a monocular depth estimator DPT [49] for this purpose [8]. Although it may not provide globally accurate depth, it allows us to ascertain the relative spatial relationships between pixels:
$\mathcal{L}_{\textit{geo}}(I_{i},D_{i})=1-\frac{\text{Cov}(D_{i},\text{DPT}(I_{i}))}{\sqrt{\text{Var}(D_{i})\text{Var}(\text{DPT}(I_{i}))}},$ (6)
@@
Figure 4: Visual Comparisons. We showcase 360 3D scene generation. In each row, from left to right, displays novel views as the camera undergoes clockwise rotation in yaw, accompanied by slight random rotations in pitch and random translations. LucidDreamer [7] hallucinates novel views from a conditioned image (indicated by a red bounding box) but lacks global semantic, stylized, and geometric consistency. In contrast, our method provides complete 360 coverage without any blind spots (black areas in baseline results), and shows globally consistent semantics.
@@
3.4 Optimization
$\mathcal{L}=\mathcal{L}_{\textit{RGB}}+\lambda_{1}\cdot\mathcal{L}_{\textit{sem}}+\lambda_{2}\cdot\mathcal{L}_{\textit{geo}}$ (7)
Here, $\mathcal{L}_{\textit{RGB}}$ represents the photometric loss on projected perspective images, which consists of $\mathcal{L}_{1}$ and a D-SSIM term as described in [25]. $\mathcal{L}_{\textit{sem}}$ denotes semantic regularization, and $\mathcal{L}_{\textit{geo}}$ signifies geometric regularization. The weights $\lambda_{1}$ and $\lambda_{2}$ are set to 0.05 each, respectively.
4 Experiments
4.1 Experiment Setting
Implementation.
Given the panorama generated from text, we resize it to $1024\times 2048$ in order to produce a dense enough point cloud by projecting per-pixel panoramic depth values into 3D space along the per-pixel camera ray directions, where the directions can be obtained according to the coordinate transformation in spherical panoramic imaging [50]. Since the panoramic depth map is predicted from the optimized geometric field, the values in local patches are guaranteed to be consistent and smooth, thereby resulting in point cloud in 3D that accurately captures the smooth geometry at the objects’ surfaces. This setup enables effective 3D Gaussians initialization within a controllable, bounded area, eliminating redundant points in empty spaces. To circumvent the common floater issue in 3DGS rendering, we disable the densification process, enhancing the overall quality and consistency of the rendered scenes.
Baseline Methods.
Our work tackles the challenging problem of unconstrained 360 generation, including both indoor and outdoor scenarios. However, the works utilizing a bounded NeRF representation using score distillation do not work very well in this case. Thus, the comparisons are conducted between DreamScene360 (ours) and the state-of-the-art LucidDreamer [7]. We use the open-source codebase of LucidDreamer, which starts from a single image and a text prompt. The framework constructs a global point cloud by progressive inpainting into 360 views and then distills a set of 3D Gaussians. In our experiments, we set the input image for the LucidDreamer to be generated from Stable Diffusion [51] v1.5 using the input text for a fair comparison.
Metrics.
4.2 Main Results
360 Scene Generation.
Comparisons with Baseline Methods.
We show visual comparisons against LucidDreamer [7] in Fig. 4. LucidDreamer involves progressive inpainting to find agreement between multiple synthesized images and tries to fuse a point cloud that can be later distilled into 3D Gaussians. Their pipeline, which inpaints each patch separately based on the same text prompt, tends to produce repetitive results especially when generating complex scenes. In comparison, our method delivers consistent results thanks to the intermediate panorama as a global 2D representation. We provide quantitative comparisons in Tab. 1, where the camera undergoes clockwise rotation in yaw, accompanied by slight random rotations in pitch and random translations to capture 4 novel views roughly representing front, back, left, and right, along with two views of up and down by 90 rotation in pitch, to mimic the exploration in an immersive 360 panoramic 3D scene.
For the bedroom text prompt, LucidDreamer begins with a view featuring a modern style bedroom with one bed but hallucinates the novel views into multiple bedrooms, while the style of the imagined bedrooms diverges from that of the starting view. In the case of the Yosemite text prompt, LucidDreamer merely replicates the waterfall seen in the initial view throughout. In conclusion, our results demonstrate global semantic, stylized, and geometric consistency, offering complete 360 coverage without any blind spots.